OpenStudy (anonymous):

If the dimensions of a rectangle are cut in half, then the new area is one fourth of the original area.true or fause

OpenStudy (anonymous):

 original area = l*b new area = l/2 * b/2 = lb/4 = original area/4 Thus, this is true!